# Decoding Quantum LDPC Codes using Collaborative Check Node Removal

### v1

@@ Abstract
The fault tolerance of quantum devices requires on-par contributions from error-correcting codes and suitable decoders. One of the most explored error-correcting codes is the family of Quantum Low-Density Parity Check (QLDPC) codes. Although faster than many of the reported decoders for QLDPC codes, iterative decoders fail due to the colossal degeneracy and short cycles intrinsic to these codes. We present a strategy to improve the performance of the iterative decoders based on a collaborative way to use the message passing of the iterative decoders and check node removal from the code’s Tanner graph. We use the concept of bit separation and generalize it to qubit separation. This gives us a metric to analyze and improve the decoder’s performance towards harmful configurations of QLDPC codes. We present a simple decoding architecture to overcome iterative decoding failures by increasing the separation of trapped qubits without incurring any significant overhead.
@@ I Introduction
Quantum Low-Density Parity Check (QLDPC) codes stand as the most suitable candidate for achieving practical fault tolerance with their low overhead [1] and good error correction properties [2, 3]. A good error-correcting code refers to code that has a constant rate and shows linear or constant distance scaling. In comparison with the classical LDPC codes, QLDPC code constructions are less trivial. The first set of construction towards acheiving the same was the hypergraph product (HGP) codes by Tillich and Zémor [3]. These QLDPC codes achieve a constant rate and a quadratic distance scaling (that is, $d\propto\sqrt{n}$) as an outcome of the tensor product between any two classical codes. The well known Toric code is a class of HGP code itself. More recent studies have presented the first constructions of asymptotically good QLDPC codes [4, 5]. For instance, Panteleev and Kalachev first proposed a generalization of the HGP construction [6], which later they renamed to Lifted Product codes [2], reporting the first QLDPC codes with almost linear distance.
@@
The performance of these QLDPC codes are far from optimal under the current decoders available. Unlike the classical LDPC codes, high degeneracy and short cycles present in the Tanner graph of the code poses a big hurdle during the decoding. Particularly iterative Belief Propagation (BP) based decoders completely fail while decoding these highly degenerate quantum codes. The state-of-the-art QLDPC decoder with a mammoth amount of success rate is the BP-OSD decoder[6]. It uses a post processing classical method based on the Ordered Statistics Decoding (OSD), to solve an inverse based problem to exhaustively find minimum weight error patterns that satisfy the syndrome. Although very accurate, it has a $\mathcal{O}(n^{3})$ time complexity in the number of physical qubits $n$ used by the error correcting code. Although, more accurate and higher order decoding of the same is too expensive for using in any practical cases. Recent attempts are made to achieve a similar accuracy with reduced time complexity [7, 8]. Also there have been certain decoders which in part, use BP and leverages from the properties of large codes, such as expansion to post-process the trapped BP outputs [9].
Different harmful configurations intrinsic to the QLDPC codes are considered as the main sorce of iterative decoding failures. Previous attempts were made based on the knowledge of these configurations and helping the decoder learn or facilitate asymmetric message passing rules to avoid failures [10, 11, 12]. In this work, we present a simple approach to assist the iterative decoder without the knowledge of these configurations. In this way we save a lot of effort towards exhaustive analysis and expensive learning based resource exploitation. We provide a notion of the qubit separation and show that removing certain mis-satisfied and symmetric check nodes (i.e. stabilizers) can improve the iterative decoder’s performance. We show that the decoder’s performance is directly linked with increased separation of the qubits belonging to the harmful configuration sets. We also leverage from the intrinsic degeneracies to estimate the separation of qubits belonging to symmetric stabilizers sets. We show a probabilistic check node removal approach like [13] offers highly separated qubits and thus can ensure a better performant decoding architecture than that of the decoders involving only iterative message passing. We observe that the proposed collaborative architecture incurs very little overhead to the overall decoding process compared to the existing post-processing methods and offers a performance bridge between iterative and the iterative + post-processing-based decoders.
@@
The paper is organized as follows: In Section II, we describe the classical and quantum error correction process and also provide necessary details of the Generalized Hypergraph Product (GHP) code that we use primarily for testing and analyzing the decoder. In Section III, we analyze the harmful configurations that arise for GHP quantum code resulting in decoding failures and discuss how one can break pass those failures by increasing the separation of the trapped qubits. Finally, we discuss the proposed collaborative decoding in Section IV and show that the proposed sub-decoding mode can indeed improve from the trapped error floor region of a BP min-sum decoder and compare the performance of the proposed work with BP and BP-OSD0 using some memory experiments for the $[[882,24]]$ and $[[1270,28]]$ GHP codes.
@@ II-A Classical Error Correction
Classical error correcting codes are a typical binary map, which encodes $k$ bits of information into $n$ bits of codeword $\mathrm{c}$, where $n>k$. The error correction mechanism can be expressed through the generator matrix $\mathbf{G}$ or the parity check matrix $\mathbf{H}$. The former method uses the fact that the rows of $\mathbf{G}$ span the set of all codewords $\mathcal{C}$. The latter defines a code through the null space of $\mathbf{H}$, that is, any codeword $\mathrm{c}\in\mathcal{C}$:
@@
This establishes the foundation to identify and correct errors. For example, if there is an error $\eta$ occurred, the codeword becomes erroneous, i.e., $\mathrm{c}\rightarrow\mathrm{y}:=\mathrm{c}+\eta$ and therefore Eq. (1) gets violated. Knowledge of the error $\eta$ is often deciphered using the syndrome $\mathrm{s}$ computed using $\mathrm{y}$ and $\mathbf{H}$ as
@@
There exist nonzero syndromes only for errors $\eta$ with a Hamming weight upper bounded by $\text{wt}(\eta)<d$, where $d$ is the minimum distance of the classical code. A decoder typically leverages the syndrome information and predicts the error occurred but the performance of the decoder is often restricted by both the decoding algorithm as well as the constraints imposed by the minimum distance of the code.
@@ II-B Quantum Error Correction
Quantum error correcting codes similar to classical codes use redundancy to encode the information of $k$ logical qubits into $n$ physical qubits. The fundamental errors that qubits suffer include continuous errors, unlike the bit flips in the classical case. A typical error in a quantum system can be expressed as a linear combination of errors in the Pauli set $\{I,X,Y,Z\}$. The Pauli-$X$ operator acts similarly to the classical bit flip error by flipping the bit-state of a qubit. Meanwhile, the Pauli-$Z$ operator changes the phase of a quantum state. The encoded logical qubit state |$\ket{\psi}_{L}$ belongs to the space of $(+1)$ eigenstates of a set of Pauli operators $\mathcal{S}\subset\mathcal{G}_{n}$, called the stabilizers of the code. Any error must anti-commute with one of the stabilizers for the decoder to be able to detect it. A one-to-one map exists between the stabilizer set and parity checks matrices for a quantum code where the rows map to the stabilizers of the code. For instance, if a quantum $[[n,k,d]]$ code has a stabilizer set $\mathcal{S}=\{s_{1},s_{2},...,s_{n-k}\}$. The parity check matrix can be defined by stacking the stabilizers in the rows to form the matrix
$\begin{bmatrix}s_{1}\\ s_{2}\\ \vdots\\ s_{n-k}\end{bmatrix}.$
The symplectic format allows a binary representation of the parity check matrix of the above representation. In this format, each of the $n$ qubit Pauli operators is represented using a $2n$ bit string. For example, one qubit Pauli operator follows
@@
The parity check matrix can be described in binary form as:
@@
This constrains the development of quantum codes. Particularly in this work, we focus on a very well-studied class of quantum codes called quantum low-density parity check (QLDPC) codes [14]. More specifically, we focus on the class of Calderbank-Shor-Steane (CSS) codes [15, 16] with a special property that the number of $1$’s in each row and column is upper bounded by a constant with the number of qubits going to infinity. This results in a very sparse parity check matrix, which is very efficient with respect to iterative decoding. The stabilizers of these codes are of the independent type, meaning they contain either Pauli-$X$ or Pauli-$Z$ operators and as a result, Eq. (6) takes a special form
@@
This ensures that decoding can be performed independently for syndromes due to bit flips and the syndromes due to phase flips. Also, Eq. (7) can now be reduced to
@@
Following this, we explore a special generalization of the HGP codes, with improved code properties that require specialized analysis.
@@ II-C Generalized Hypergraph Product (GHP) Codes
An intuitive way to construct QLDPC codes based on Eq. (9), is using two commuting matrices and stack them together to form the parity check matrix of the QLDPC code [17]. Panteleev and Kalachev use this intuition to propose a generalization of the standard hypergraphproduct codes [6]. In their work they consider two square matrices, elements of which belongs to a ring called ring of circulants defined below in II.1.
@@
A ring of circulant $\mathbb{A}_{L}$ satisfies all the properties of a ring with its members being the circulant permutation matrices or sometimes referred to as the polynomials. The whole ring is characterized by a parameter called "lift", which is essentially the dimension of the permutation matrices.
@@
where A is a matrix $A=[a_{ij}]_{m\times n}$, such that $a_{ij}\in\mathbb{A}_{L}$ and $m=n$. Further, $b$ is a polynomial in $\mathbb{A}_{L}$ and $I_{m}$ and $I_{n}$ are diagonal matrices with its diagonal entries being the identity of the ring. For example, we will extensively use the $[[882,24]]$ code discussed in [6], which uses
@@
where $x^{i},\,\,\forall i\in\{1,...,L-1\}$ are the polynomials with it’s columns shifted by $i$ from the identity. Also the lift for the ring is set as $L=63$, thus all this polynomials are $(63\times 63)$ matrices over $\mathbb{F}_{2}$.
@@ II-D Iterative Decoding
We discussed in the previous sections that errors anti-commuting with the stabilizers result in non-zero syndrome and indicate the occurrence of errors. Based on the received syndrome $\mathrm{s}$, the decoding problem is a search for the error, which maximizes the posterior probability:
@@
where $E$ is the coset of an error $e\in\mathcal{G}_{n}$ such that $E:=e\mathcal{S}$. $\mathcal{E}$ is a collection of all such possible cosets. The estimated error $\hat{e}$ of Eq. (14) can be any operator from the maximum probability coset $E$, as members of a single coset has the same logical effect on the state.
Throughout this article we use iterative decoding and BP decoding in an equivalent manner. BP is an iterative message-passing algorithm due to its continuous back-and-forth information-passing characteristics over a bipartite Tanner graph of the code. BP primarily evaluates the marginals of Eq. (13), i.e.
@@
Therefore it so happens that the maximized $\hat{e}$ in (13) does not always satisfy Eq. (14). Many variants of the BP algorithm exist, primarily depending on the application of the algorithm. We describe very briefly the min-sum algorithm, which we used for the numerical experiments and also offers better numerical stability than the others. This algorithm performs series of binary message passing from data qubit nodes to it’s adjacent check nodes and vice-versa, producing soft outputs for each data qubit nodes. We use $u_{i}$ and $v_{j}$ to denote the $i^{th}$ data and $j^{th}$ check nodes respectively. Also, we assume the notation $m^{k}_{a\rightarrow b}$ to portray message passed at the $k^{th}$ iteration of the algorithm from node $a$ to node $b$ on the Tanner graph. Following we describe an overall brief summary of the standard message passing rule of the min-sum BP:
@@
    Initialization: Initially from each data qubits a message $m^{0}_{u_{i}\rightarrow v_{j}}$ is passed as follows:
    $m^{0}_{u_{i}\rightarrow v_{j}}=\lambda_{i}=\log(\frac{1-p_{i}}{p_{i}});\,\,\forall i,j:\mathbf{H}_{ji}=1.$ (16)
@@
    Check to data qubit messages: For the subsequent iterations, message $m^{k}_{v_{j}\rightarrow u_{i}}$ is passed from each check node to data qubits. Although note we exclude the message passing to the qubit node, which was immediately involved in message passing at ${k-1}^{th}$ iteration. We define a quantity $w=\min_{u_{p}:\,N(v_{j})/u_{i}}\{|m^{k-1}_{u_{p}\rightarrow v_{j}}|\}$ and the message passed is as follows:
    $m^{k}_{v_{j}\rightarrow u_{i}}=(-1)^{\mathrm{s}_{j}}\alpha\left(\prod_{u_{p}:\,N(v_{j})\backslash u_{i}}sign(m^{k-1}_{u_{p}\rightarrow v_{j}})\right)w,$ (17)
    where, $\alpha$ is a scaling factor and can affect the convergence and $\mathrm{s}_{j}$ is the syndrome value corresponding to the check node $v_{j}$.
@@
    Data qubit to check messages: Except the initialization as described, at each iteration $k$, the messages passed from data qubits to check nodes follows:
    $m^{k}_{u_{i}\rightarrow v_{j}}=\lambda_{i}+\sum_{v_{p}:N(u_{i})\backslash v_{j}}m^{k-1}_{v_{p}\rightarrow u_{i}}$ (18)
@@
    Hard decision: At the end of each iteration, a soft output is evaluated for each qubit node. The soft output represents the collective response received from all the check nodes and is as follows:
    $\gamma_{i}\leftarrow\lambda_{i}+\sum_{v_{p}:N(u_{i})}m^{k}_{v_{p}\rightarrow u_{i}}.$ (19)
@@
The whole process continues until all the unsatisfied checks are satisfied or a pre-determined number of message-passing iterations is reached, and the syndrome is still non-zero. In the former, the decoding is said to have succeeded, and the latter reflects a decoding failure. A cycle-free Tanner graph guarantees the success of BP within a finite number of iterations, although QEC offers plenty of short cycles along with inevitable degeneracies, which makes BP fail miserably. In the following section, we discuss one such case where some inherent configurations of the GHP code render BP an ineffective decoder.
@@
III Trapping Sets and Separation Vectors of GHP codes
Iterative message passing-based decoders are vulnerable when it comes to certain error configurations. These are heavily studied and formally named as the Trapping sets (TS) and we will use $\mathcal{T}_{s}$ to denote the set of data qubits that are trapped. If $\mathcal{T}_{s}$ contains $a$ data qubits and $b$ odd degree stabilizer checks then it is a $(a,b)$ trapping set of the code. We use the notation $V_{1}$ to describe the set of qubits $v\in\mathcal{T}_{s}$, such that $v$ is adjacent to odd degree stabilizer checks. Note that there can be many possible configurations for any $(a,b)$ TS, depending on the geometry of the code and the decoder under consideration.

### v2

@@ Abstract
Fault tolerance of quantum protocols require on-par contributions from error-correcting codes and it’s suitable decoders. One of the most explored error-correcting codes is the family of Quantum Low-Density Parity Check (QLDPC) codes. Although faster than many of the reported decoders for QLDPC codes, iterative decoders fails to produce suitable success rates due to the colossal degeneracy and short cycles intrinsic to these codes. We present a strategy to improve the performance of the iterative decoders based on a collaborative way to use the message passing of the iterative decoders and stabilizer check node removal from the quantum code’s Tanner graph. We particularly introduce a notion of “qubit separation", which gives us a metric to analyze and improve the min-sum Belief Propagation (BP) based iterative decoder’s performance towards harmful configurations of QLDPC codes. We further show that an integration of information measurements (IM) for qubits and it’s adjacent stabilizer checks, can be exploited to extract far better performing results from the collaborative decoding architecture compared to it’s classical predecessor. We analyze the performance of the proposed collaborative decoding architecture, in the context of Generalized Hypergraph Product (GHP) codes. We discuss that the collaborative decoding architecture overcomes iterative decoding failures regarding the harmful trapping set configurations by increasing the separation of trapped qubits without incurring any significant overhead.
Index Terms:
Quantum Error Correction, QLDPC codes, Decoders, Belief Propagation, Trapping sets, Qubit Separation, Generalized Hypergraph Product codes.
@@ I Introduction
Quantum Error correction has been a primary focus of many to realize the quantum protocols with high fidelity. Over the years, Quantum Low-Density Parity Check (QLDPC) codes have emerged as the most suitable candidate for achieving practical fault tolerance due to their low overhead [1] and good error correction properties [2, 3]. A good error-correcting code refers to code that has a constant rate and shows linear or constant distance scaling. In comparison with the classical LDPC codes, good QLDPC code constructions are less trivial. The first set of construction towards achieving the same was the Hypergraph Product (HGP) codes by Tillich and Zémor [3]. These QLDPC codes achieve a constant rate and a quadratic distance scaling (that is, $d\propto\sqrt{n}$) as an outcome of the tensor product between any two classical codes. The well-known Toric code is a class of HGP code itself [4]. More recent studies have presented the first constructions of asymptotically good QLDPC codes [5, 6]. For instance, Panteleev and Kalachev proposed a generalization of the HGP construction [7], which later they renamed to Lifted Product codes [2], reporting the first QLDPC codes with almost linear distance.
@@
The performance of these QLDPC codes is far from optimal under the current available decoders. Unlike the classical LDPC codes, the high degeneracy and short cycles present in the Tanner graph of the code pose a big hurdle during decoding. Particularly iterative Belief Propagation (BP) based decoders completely fail while decoding these highly degenerate quantum codes. For instance, the surface code exhibits no threshold under the BP decoding. In this article, we use both the iterative and BP decoders alternately on an equal footing to represent the same class of message-passing-based decoder. The state-of-the-art QLDPC decoder, which has the capability of improving the decoding success with a mammoth leap is the BP+OSD decoder [7]. This particular decoder uses a post-processing classical method based on the Ordered Statistics Decoding (OSD), to solve a matrix inversion problem and would exhaustively find minimum weight error patterns that satisfy the syndrome. Although very accurate, even zeroth-order OSD post-processing has a $\mathcal{O}(n^{3})$ time complexity in the number of physical qubits $n$ used by the error correcting code. Further improvements in the accuracy require higher-order post-processing of the same and are too expensive to use in any practical cases. Recent attempts have been made to achieve similar accuracy with reduced time complexity [8, 9]. In addition, certain decoders have been developed thatat use BP anleverage properties of large codes, such as expansion to post-process the trapped BP outputs [10].
Different harmful configurations, popularly known as Trapping Sets (TS), intrinsic to the QLDPC codes, are considered the main source of iterative decoding failures. Attempts are made based on the knowledge of these configurations to assist the decoder learn or facilitate asymmetric message passing rules to avoid failures [11, 12, 13, 14]. In this work, we present a simple approach to assist the iterative decoder without any active monitoring requirement for the knowledge of these configurations in decoding. Our motivation is to inhibit a set of information flows that might result in convergence failures for any iteration of the message passing iterative decoder. We introduce the notion of qubit separation and evaluate information measurement (IM) values to identify the high risk stabilizer check nodes for a given set of violated stabilizer check nodes, i.e. syndrome. We propose that these high-risk check nodes having high IM values limit the separation of certain data qubits adjacent to it and show that the iterative decoder is unable to converge for this set of data qubits. Therefore, we conclude that removing these high-risk check nodes (i.e. stabilizers) can improve the iterative decoder’s performance, and doing so, we establish that the decoder’s performance is directly linked to increased separation of the data qubits belonging to the harmful configurations.
A similar intuition was adopted in [15], where the authors removed a set of data qubits in each iteration of the BP decoder as a post-processing step. This essentially deactivates some stabilizer check node in each BP iteration, and they execute such deactivation according to the net soft information of the adjacent data qubits. The key difference between our work and [15] is that we do not modify the codespace, i.e. our decoding sub-routines directly remove stabilizer check nodes without any straight exclusion of information passed from the data qubits to other check nodes. Also, we will see that our proposed decoder turns off multiple stabilizer check nodes in a single iteration, whereas the stabilizer inactivation in [15] is done one by one for each iteration of BP. This, on average, converges the stuck message passing BP quickly, which is further supported by the additional sub-routine of information measurement and quantified through improvement of qubit separation. The proposed way of improving the iterative decoder’s performance saves a lot of effort towards exhaustive analysis and expensive learning based resource exploitation. We also discuss improvements over the probabilistic check node removal approach of Kang et al. [16] and discuss how the proposed algorithmic sub-routines ensure highly separated qubits and thus can offer a better-performing decoding architecture than that of the decoders involving only iterative message passing. We observe that the proposed collaborative architecture incurs very little overhead to the overall decoding process compared to the existing post-processing methods and still offers significant success rate improvements compared to standard iterative decoders like min-sum BP.
@@
The paper is organized as follows: In Section II, we describe the classical and quantum error correction process and also provide necessary details of the Generalized Hypergraph Product (GHP) code that we use primarily for testing and analyzing the decoder. In Section III, we analyze the harmful configurations that arise for GHP quantum code resulting in decoding failures and discuss how one can break pass those failures by increasing the separation of the trapped qubits. Finally, we discuss the proposed collaborative decoding in Section IV and show that the proposed sub-decoding mode can indeed improve from the trapped error floor region of a min-sum BP decoder and compare the performance of the proposed work with BP and BP+OSD00 using memory experiments for the $[[882,24]]$ and $[[1270,28]]$ GHP codes. Additionally, we also provide proof of existing thresholds for the Generalized Bicycle (GB) codes with the use of the proposed decoder.
@@ II-A Classical Error Correction
Classical error correcting codes are a typical binary map, which encodes $k$ bits of information into $n$ bits of codeword $\mathrm{c}$, where $n>k$. The error correction mechanism can be expressed through the generator matrix $\mathbf{G}$ or the parity check matrix $\mathbf{H}$. The former method uses the fact that the rows of $\mathbf{G}$ span the set of all codewords $\mathcal{C}$. The latter defines a code through the null space of $\mathbf{H}$, that is, any codeword $\mathrm{c}\in\mathcal{C}$ satisfy:
@@
This establishes the foundation to identify and correct errors. For example, if there is an error $\eta$ occurred, the codeword becomes erroneous, i.e., $\mathrm{c}\rightarrow\mathrm{y}:=\mathrm{c}+\eta$ and therefore Eq. (1) gets violated. Knowledge of the error $\eta$ is often deciphered using the syndrome $\mathrm{s}$, calculated using $\mathrm{y}$ and $\mathbf{H}$ as
@@
There exist non-zero syndromes only for errors $\eta$ with a Hamming weight upper bounded by $\text{wt}(\eta)<d$, where $d$ is the minimum distance of the classical code. A decoder typically leverages the syndrome information and predicts the error that occurred, but the performance of the decoder is often restricted by both the decoding algorithm and the constraints imposed by the minimum distance of the code.
@@ II-B Quantum Error Correction
Quantum error correcting codes, similar to classical codes, use redundancy to encode the information of $k$ logical qubits into $n$ physical qubits. The fundamental errors that qubits suffer include a continuous set of errors, unlike the classical case, where the only source of errors is bit flips at discrete positions. A typical error in a quantum system can be expressed as a linear combination of errors in the Pauli set $\{I,X,Y,Z\}$. The Pauli-$X$ operator acts similarly to the classical bit flip error by flipping the state of a qubit. Meanwhile, the Pauli-$Z$ operator changes the phase of a quantum state. The encoded logical qubit state |$\ket{\psi}_{L}$ belongs to the space of $(+1)$ eigenstates of a set of Pauli operators $\mathcal{S}\subset\mathcal{G}_{n}$, called the stabilizers of the code. Any error must anti-commute with one of the stabilizers for the decoder to be able to detect it. A one-to-one map exists between the stabilizer set and the parity check matrix (PCM) of a quantum code. Each row of the PCM maps to a stabilizer of the code. For instance, if a quantum $[[n,k,d]]$ code has a stabilizer set $\mathcal{S}=\{s_{1},s_{2},...,s_{n-k}\}$. The parity check matrix can be defined by stacking the stabilizers in the rows of a matrix of the form:
$\mathbf{H}=\begin{bmatrix}s_{1}\\ s_{2}\\ \vdots\\ s_{n-k}\end{bmatrix}.$
The symplectic format allows a binary representation of the above parity check matrix. In this format, each of the $n$ qubit Pauli operators is represented using a $2n$ bit string. For example, one qubit Pauli operator follows;
@@
Therefore, the parity check matrix can be described in binary form as:
@@
This constrains the development of quantum codes. Particularly in this work, we focus on a very well-studied class of quantum codes called quantum low-density parity check (QLDPC) codes [17]. More specifically, we focus on the class of Calderbank-Shor-Steane (CSS) codes [18, 19] with a special property that the number of $1$’s in each row and column is upper bounded by a constant with the number of qubits going to infinity. This results in a very sparse parity check matrix, which is very efficient with respect to iterative decoding. The stabilizers of these codes are of the independent type, meaning they contain either Pauli-$X$ or Pauli-$Z$ operators, and as a result, Eq. (6) takes a special form
@@
This ensures that the decoding can be performed independently for syndromes due to bit flips and the syndromes due to phase flips. Also, Eq. (7) can now be reduced to
@@
Following this, we explore a very special class of QLDPC codes, namely the Generalized Hypergraph Product (GHP) codes. It is the generalization of the hypergraph construction, and holds great popularity due to its improved asymptotic performance and better distance scaling.
@@ II-C Generalized Hypergraph Product (GHP) Codes
An intuitive way to construct QLDPC codes based on Eq. (9), is using two commuting matrices and stack them together to form the parity check matrix of the QLDPC code [20]. Panteleev and Kalachev use this intuition to propose a generalization of the standard hypergraph product codes [7]. In their work, they consider two square matrices, elements of which belong to a ring called the ring of circulants defined below in II.1.
@@
A ring of circulant $\mathbb{A}_{L}$ satisfies all the properties of a ring, with its members being the circulant permutation matrices, or sometimes referred to as the polynomials. The whole ring is characterized by a parameter called "lift", which is essentially the dimension of the square permutation matrices.
@@
where A is a matrix $A=[a_{ij}]_{m\times n}$, such that $a_{ij}\in\mathbb{A}_{L}$ and $m=n$. Further, $b$ is a polynomial in $\mathbb{A}_{L}$ and $I_{m}$ and $I_{n}$ are diagonal matrices with their diagonal entries being the identity of the ring. For example, we will extensively use the $[[882,24]]$ code discussed in [7], which uses
@@
where $x^{i},\,\,\forall i\in\{1,...,L-1\}$ are the polynomials with it’s columns shifted by $i$ positions, compared to the corresponding identity element. Also, the lift for the ring is set to $L=63$; thus, all these polynomials are $(63\times 63)$ matrices over $\mathbb{F}_{2}$. We also experiment with the $[[1270,28]]$ GHP code (see the appendix of [7] for details).
@@ II-D Iterative Decoding
We discussed in the previous sections that errors anti-commuting with the stabilizers result in a non-zero syndrome, indicating the occurrence of errors. Based on the received syndrome $\mathrm{s}$, the decoding problem is a search for the error, which maximizes the posterior probability:
@@
where $E$ is the coset of an error $e\in\mathcal{G}_{n}$ such that $E:=e\mathcal{S}$. $\mathcal{E}$ is a collection of all such possible cosets. The estimated error $\hat{e}$ of Eq. (14) can be any operator from the maximum probability coset $E$, as members of a single coset have the same logical effect on the state.
BP is an iterative message-passing algorithm due to its continuous back-and-forth information-passing characteristics over a bipartite Tanner graph of the code. BP primarily evaluates the marginals of Eq. (13), i.e.
@@
Therefore, it so happens that the maximized $\hat{e}$ in (13) does not always satisfy Eq. (14). Many variants of the BP algorithm exist, primarily depending on the application of the algorithm. We describe very briefly the min-sum algorithm, which we used for the numerical experiments, and also offers better numerical stability than the others. This algorithm performs a series of binary message passing from data qubit nodes to its adjacent check nodes and vice-versa, producing soft outputs for each data qubit node. We use $v_{i}$ and $c_{j}$ to denote the $i^{th}$ data and $j^{th}$ check nodes respectively. Also, we assume the notation $m^{k}_{a\rightarrow b}$ to portray the message passed at the $k^{th}$ iteration of the algorithm from node $a$ to node $b$ on the Tanner graph. Following we describe an overall brief summary of the standard message passing rule of the min-sum BP:
@@
    Initialization: Initially from each data qubits a message $m^{0}_{v_{i}\rightarrow c_{j}}$ is passed as follows:
    $m^{0}_{v_{i}\rightarrow c_{j}}=\lambda_{i}=\log(\frac{1-p_{i}}{p_{i}});\,\,\forall i,j:\mathbf{H}_{ji}=1.$ (16)
@@
    Check to data qubit messages: For the subsequent iterations, message $m^{k}_{c_{j}\rightarrow v_{i}}$ is passed from each check node to data qubits. Although note we exclude the message passing to the qubit node, which was immediately involved in the same message passing route at ${k-1}^{th}$ iteration. We define a quantity $w=\min_{v_{p}:\,N(c_{j})/v_{i}}\{|m^{k-1}_{v_{p}\rightarrow c_{j}}|\}$ and the message passed is as follows:
    $m^{k}_{c_{j}\rightarrow v_{i}}=(-1)^{\mathrm{s}_{j}}\alpha\left(\prod_{v_{p}:\,N(c_{j})\backslash v_{i}}\mathrm{sign}(m^{k-1}_{v_{p}\rightarrow c_{j}})\right)w,$ (17)
    where, $\alpha$ is a scaling factor and can affect the convergence and $\mathrm{s}_{j}$ is the syndrome value corresponding to the check node $c_{j}$. Also, here and in the rest of the article, we denote the set of neighborhoods of any data qubit $q_{i}$ or stabilizer check $c_{j}$ as $N(q_{i})$, $N(c_{j})$, respectively.
@@
    Data qubit to check messages: Except for the initialization as described, at each iteration $k$, the messages passed from data qubits to check nodes follows:
    $m^{k}_{v_{i}\rightarrow c_{j}}=\lambda_{i}+\sum_{c_{p}:N(v_{i})\backslash c_{j}}m^{k-1}_{c_{p}\rightarrow v_{i}}$ (18)
@@
    Hard decision: At the end of each iteration, a soft output is evaluated for each qubit node. The soft output represents the collective information received from all the check nodes and is as follows:
    $\gamma_{i}\leftarrow\lambda_{i}+\sum_{c_{p}:N(v_{i})}m^{k}_{c_{p}\rightarrow v_{i}}.$ (19)
@@
The whole process continues until all the unsatisfied checks are satisfied or a pre-determined number of message-passing iterations have been completed, yet the number of unsatisfied checks is still non-zero. In the former, the decoding is said to have succeeded, and the latter reflects a decoding failure. A cycle-free Tanner graph guarantees the success of BP within a finite number of iterations, although QEC offers plenty of short cycles along with inevitable degeneracies, which makes BP fail miserably. In the following section, we discuss such cases where some inherent configurations of the GHP code renders min-sum BP as an ineffective decoder.
@@
III Trapping Sets and Qubit Separation of QLDPC codes
Iterative message-passing-based decoders are vulnerable when it comes to certain error configurations. These are heavily studied and formally named as the Trapping sets (TS). We use $\mathcal{T}_{s}$ to denote the set of data qubits that are trapped in any TS. If $\mathcal{T}_{s}$ contains $a$ data qubits and has $b$ number of odd degree stabilizer checks adjacent to $\mathcal{T}_{s}$, then it is a $(a,b)$ TS of the code. Any $(a,b)$ TS has many possible configurations, depending on the code structure and the decoder under consideration. In the following, we primarily discuss the necessary concepts of TS in the context of GHP codes and show the typical structure of the GHP code trapping sets. Later, we discuss how the notion of qubit separation is exploited to identify the stabilizer check nodes, whose contributions, if ignored, both the trapped qubit’s separation and the iterative decoder’s performance are increased.